# Specific heat of Gd3+- and Eu2+-based magnetic compounds

### v1

@@ VIII Appendix
Here we calculate the mean value and fluctuations of a random variable $x$ with different intensities. For simplicity we consider only two intensities $J_{1}$ and $J_{2}$. Extension to the general case is straightforward. We consider $z_{1}$ attempts in which the probability of the result $J_{1}$ is $p$ and that of $-J_{1}$ is $q=1-p$. In addition, there are $z_{2}$ attempts with probability $r$ of finding the result $J_{2}$ and $s=1-r$ for the result $-J_{2}$. At the end we will take $r=p$. For non-frustrated systems, both $J_{\delta}$ can be chosen positive. Clearly, the expectation value of $x$ is
@@
On the other hand, considering the probability of each individual event one has
$\langle x^{2}\rangle=\sum\limits_{n=0}^{z_{1}}\sum\limits_{m=0}^{z_{2}}\binom{z_{1}}{n}\binom{z_{2}}{m}p^{n}q^{z_{1}-n}r^{m}s^{z_{2}-m}\left[nJ_{1}-(z_{1}-n)J_{1}+mJ_{2}-(z_{2}-m)J_{2}\right]^{2}.$ (8)
@@
$O=\left[J_{1}\left(p\frac{\partial}{\partial p}-q\frac{\partial}{\partial q}\right)+J_{2}\left(r\frac{\partial}{\partial r}-s\frac{\partial}{\partial s}\right)\right],\text{ and }(p+q)^{z_{1}}=\sum\limits_{n=0}^{z_{1}}\binom{z_{1}}{n}p^{n}q^{z_{1}-n},$ (9)
and a similar expression for $(r+s)^{z_{2}}$, in which $p$ and $q$ as well as $r$ and $s$ are taken as independent variables in the derivatives, one realizes that the expectation value Eq. (8) can be written as
@@
Performing the calculation we find after some algebra
@@
$\frac{\langle x^{2}\rangle-\langle x\rangle^{2}}{\langle x\rangle^{2}}=\frac{4p(1-p)}{2p-1}\frac{z_{1}J_{1}^{2}+z_{2}J_{2}^{2}}{\left(z_{1}J_{1}+z_{2}J_{2}\right)^{2}}$ (12)

### v2

@@ VIII Appendix
Here we calculate the mean value and fluctuations of a random variable $x$ with different intensities  [^1] The same result can be obtained considering each bond as an independent variable.. For simplicity we consider only two intensities $J_{1}$ and $J_{2}$. Extension to the general case is straightforward. We consider $z_{1}$ attempts in which the probability of the result $J_{1}$ is $p$ and that of $-J_{1}$ is $q=1-p$. In addition, there are $z_{2}$ attempts with probability $r$ of finding the result $J_{2}$ and $s=1-r$ for the result $-J_{2}$. At the end we will take $r=p$. For non-frustrated systems, both $J_{\delta}$ can be chosen positive. Clearly, the expectation value of $x$ is
@@
On the other hand, considering the probability of each individual event, we have:
$\langle x^{2}\rangle$ $=$ $\sum\limits_{n=0}^{z_{1}}\sum\limits_{m=0}^{z_{2}}\binom{z_{1}}{n}\binom{z_{2}}{m}p^{n}q^{z_{1}-n}r^{m}s^{z_{2}-m}$ (10)
×$\times\left[nJ_{1}-(z_{1}-n)J_{1}+mJ_{2}-(z_{2}-m)J_{2}\right]^{2}.$
@@
$O=\left[J_{1}\left(p\frac{\partial}{\partial p}-q\frac{\partial}{\partial q}\right)+J_{2}\left(r\frac{\partial}{\partial r}-s\frac{\partial}{\partial s}\right)\right],$
[…]
and a similar expression for $(r+s)^{z_{2}}$, where $p$ and $q$ as well as $r$ and $s$ are taken as independent variables in the derivatives, one realizes that the expectation value Eq. (10) can be written as
@@
Performing the calculation, we find after some algebra
@@
$\frac{\langle x^{2}\rangle-\langle x\rangle^{2}}{\langle x\rangle^{2}}=\frac{4p(1-p)}{(2p-1)^{2}}\frac{z_{1}J_{1}^{2}+z_{2}J_{2}^{2}}{\left(z_{1}J_{1}+z_{2}J_{2}\right)^{2}}.$ (14)
Naturally, if $\langle x\rangle=0$ as it is the case for temperatures above the critical temperature $T_{C}$, Eq. (14) gives a divergent result. In this case a meaningful comparison should involve the fluctuations above $T_{C}$ and the value of $\langle x\rangle=0$ at some point below $T_{C}$.